OpenStudy (anonymous):

find the area of a rectangle heating duct grill having a length of 2 yds and a width of 3/4yd.

OpenStudy (nikato):

Area= length x width

OpenStudy (anonymous):

I did I got 3/2 but its wrong

OpenStudy (nikato):

Is there a picture to it?

OpenStudy (anonymous):

no its aword problem

OpenStudy (nikato):

Is that all there is?

OpenStudy (anonymous):

it just says type a whole number or a mixed number that's it

OpenStudy (nikato):

Mixed number? 3/2 is an improper fraction!

OpenStudy (anonymous):

well I am not good at math

OpenStudy (nikato):

It'll be 1 1/2

OpenStudy (anonymous):

oh okay thank u

OpenStudy (nikato):

3/2 is an improper fraction

OpenStudy (nikato):

Ur welcome

jimthompson5910 (jim_thompson5910):

Divide 3/2 using long division or a calculator to get 1.5 The whole part is 1 -------------------------------------------------- Since the decimal portion is 0.5, you multiply it by the denominator 2 to get 0.5*2 = 1 The remainder is 1 -------------------------------------------------- So these two facts mean \(\Large \frac{3}{2} = 1\frac{1}{2}\)